# Whistle: Data-Efficient Multilingual and Crosslingual Speech Recognition via Weakly Phonetic Supervision

### v1

@@ VII Conclusions and future work
This paper starts from examining the pros and cons of the three main approaches for MCL-ASR - supervised pre-training with phonetic transcription or graphemic transcription, and self-supervised pre-training. We find that pre-training with phonetic supervision has been underappreciated so far for MCL-ASR, while conceptually it is more advantageous for information sharing between different languages. This paper explores the approach of pre-training with weakly phonetic supervision towards data-efficient MCL-ASR, which is called Whistle. We relax the requirement of gold-standard human-validated phonetic transcripts, and obtain IPA based transcripts by leveraging Phonetisaurus (an FST based G2P toolkit) with LanguageNet G2P FSTs. We construct a common experimental setup based on the CommonVoice dataset, called CV-Lang10, with 10 seen languages and 2 unseen languages (Polish and Indonesian). A set of experiments are conducted on CV-Lang10 to compare, as fair as possible, the three approaches under the common setup for MCL-ASR. Training with weakly phonetic supervision (though somewhat noisy) and decoding with PROLEXs, with phonemes serving as an interface between acoustics and text, is found to obtain superior results in MCL-ASR in our experiments, in terms of speech recognition for seen languages, crosslingual performance for unseen languages with different amounts of few-shot data, overcoming catastrophic forgetting, and training efficiency. Moreover, phoneme-based models naturally overcome language imbalance and can be efficiently trained on natural data mixing, while subword-based models need careful tokenization and data mixing in training. When training data is more limited, phoneme supervision can achieve better results compared to subword supervision and self-supervision, thereby providing higher data-efficiency.
This work demonstrates some advantages of weakly phonetic supervision towards data-efficient MCL-ASR. There are interesting directions for future work. First, we preliminarily sidestep the problem how tones should be incorporated in pre-training multilingual phoneme-based models, since the 12 languages examined in this paper are all non-tonal languages. There have been some effort towards addressing this problem [55]. Second, this work mainly uses WFST based decoding with PROLEXs. Better methods of decoding from phonemes could be explored in future, such as based on sequence-to-sequence models [59]. Third, scaling the approach of Whistle with more languages and more data is expected to achieve increasingly better MCL-ASR performance. Meanwhile, it is worthwhile to investigate how to incrementally learn from new languages with a non-stationary stream Continual learning methods such as based on prompt pool [60, 61] could be incorporated into MCL-ASR.

### v2

@@ VII Conclusions and future work
This paper starts from examining the pros and cons of the three main approaches for MCL-ASR - supervised pretraining with phonetic transcription or graphemic transcription, and self-supervised pretraining. We find that pretraining with phonetic supervision has been underappreciated so far for MCL-ASR, while conceptually it is more advantageous for information sharing between different languages. This paper explores the approach of pretraining with weakly phonetic supervision towards data-efficient MCL-ASR, which is called Whistle. We relax the requirement of gold-standard human-validated phonetic transcripts, and obtain IPA based transcripts by leveraging Phonetisaurus (an FST based G2P toolkit) with LanguageNet G2P FSTs. We construct a common experimental setup based on the CommonVoice dataset, called CV-Lang10, with 10 seen languages and 2 unseen languages (Polish and Indonesian). A set of experiments are conducted on CV-Lang10 to compare, as fair as possible, the three approaches under the common setup for MCL-ASR. Training with weakly phonetic supervision (though somewhat noisy) and decoding with PROLEXs, with phonemes serving as an interface between acoustics and text, is found to obtain superior results in MCL-ASR in our experiments, in terms of speech recognition for seen languages, crosslingual performance for unseen languages with different amounts of few-shot data, overcoming catastrophic forgetting, and training efficiency. Moreover, phoneme-based models naturally overcome language imbalance and can be efficiently trained on natural data mixing, while subword-based models need careful tokenization and data mixing in training. When training data is more limited, phoneme supervision can achieve better results compared to subword supervision and self-supervision, thereby providing higher data-efficiency.
This work demonstrates some advantages of weakly phonetic supervision towards data-efficient MCL-ASR. There are interesting directions for future work. First, we preliminarily sidestep the problem how tones should be incorporated in pretraining multilingual phoneme-based models, since the 12 languages examined in this paper are all non-tonal languages. There have been some effort towards addressing this problem [60]. Second, this work mainly uses WFST based decoding with PROLEXs. Better methods of decoding from phonemes could be explored in future, such as based on sequence-to-sequence models [64]. Third, scaling the approach of Whistle with more languages and more data is expected to achieve increasingly better MCL-ASR performance. Meanwhile, it is worthwhile to investigate how to incrementally learn from new languages with a non-stationary stream Continual learning methods such as based on prompt pool [65, 66] could be incorporated into MCL-ASR.